# SeamlessEdit: Background Noise Aware Zero-Shot Speech Editing with in-Context Enhancement

### v1

@@ Abstract
With the fast development of zero-shot text-to-speech technologies, it is possible to generate high-quality speech signals that are indistinguishable from the real ones. Speech editing, including speech insertion and replacement, appeals to researchers due to its potential applications. However, existing studies only considered clean speech scenarios. In real-world applications, the existence of environmental noise could significantly degrade the quality of the generation. In this study, we propose a noise-resilient speech editing framework, SeamlessEdit, for noisy speech editing. SeamlessEdit adopts a frequency-band-aware noise suppression module and an in-content refinement strategy. It can well address the scenario where the frequency bands of voice and background noise are not separated. The proposed SeamlessEdit framework outperforms state-of-the-art approaches in multiple quantitative and qualitative evaluations.
keywords:
speech editing, noisy speech, in-context learning, speech synthesis
@@ 1 Introduction
Several techniques can effectively improve the quality of audio generation and eliminate the acoustic discrepancy in the segmented boundaries. Typically, fade-in and fade-out techniques are used to smooth the transitions between cut-and-paste audio segments [4]. However, intricate editing procedures may generate artifacts and degrade audio quality. Text prompts are viable for generating the speech signal following a specific style  [5]. P-Flow in  [6] uses a prompt-based speaker-conditioned text representation to enhance quality. In addition, context-aware methods have been proposed to eliminate unnatural edited boundaries  [7]. FluentSpeech [8] alleviates the over-smoothing problem using iterative refinement for stutter removal. Mapache [9] addresses the problem of left-to-right generation in autoregressive (AR) models by a non-AR parallel transformer. FluentEdit [10] applies the acoustic and prosody consistency loss when training the diffusion model for text-based editing. InstructSpeech [11] is a multitask model and applies a large language model for training.
@@
This study proposes a noise-resilient speech editing framework, SeamlessEdit, that performs in-context refinement using coarsely separated speech and fine-grained noise-suppressed speech for seamless speech-noise integration. The framework extracts the coarsely separated speech and the background noises using a score-based diffusion model. Then, we leverage a sparse Bayesian learning (SBL) filtering approach that preserves major sparse voice components and suppresses residual noise. This noise-suppressed speech is incorporated into the decoder for in-context refinement first introduced to the speech editing field, enabling the enhancement of auditory quality and seamless boundary transitions. The experiments on the EARS-WHAM dataset [16] using our proposed SeamlessEdit demonstrate 29.01% NMOS improvement and 24.28% SMOS improvement compared to the state-of-the-art speech editing approaches. The results of human evaluation also reveal significantly improved performance of the proposed model in generating indistinguishable synthetic audio that effectively blends realistic background noises, original speech, and edited speech. Qualitative and quantitative analyses are also conducted to show the efficacy of the proposed model.
The demo page and code will be released after paper acceptance. The detail of demo can be seen from the Supplementary Document.
@@ 2 Methods
Figure 1: The proposed SeamlessEdit framework separates human voice $X_{s}$ and suppresses residual noise to derive an edited speech $X_{le}$. An in-context refinement enhances the editing of $X_{s}$ using $X_{le}$ for indistinguishable noisy editing results. The SBL filter is also adopted to improve the robustness to noise.
@@
For the experiments in this work, we utilize the EARS-WHAM dataset that combines the recently proposed EARS (expressive anechoic recordings of speech) [16] dataset and the diverse background noises from WHAM! (WSJ0 Hipster Ambient Mixtures) dataset [17]. EARS comprises high-quality anechoic 16kHz speech from diverse speakers. It consists of a variety of emotional, narrative, non-verbal, and conversational speech signals. We evaluate speech editing performance on the EARS-WHAM dataset and compare the results on clean and noisy subsets. We use 886 samples introduced in [16] paired with 594 background noises from the WHAM! dataset across 198 SNR levels. The speech signals have a mean length of 15.58 seconds and contain 36.3 spoken words on average.
@@ 2.1 SeamlessEdit: Noisy Speech Editing Framework
Given an input speech sequence $X=\{X_{1},...,X_{T}\}$ with length $T$, the $m^{th}$ to $(m+k)^{th}$ samples specify the targeted editing region, and thus the edited speech is denoted as $X_{e}=\{X_{1},...,X_{m},...,X_{m+k}\,...,X_{T}\}$. The proposed SeamlessEdit framework consists of a speech separation module (§$\S$2.2) followed by a frequency-aware noise suppression module (§$\S$2.3) to extract the separated signal $X_{s}$, the noise suppressed signal $X_{l}$, and the background signal $X_{n}$. The proposed in-context refinement strategy enhances the pre-trained editing model (§$\S$2.4) to generate edited voices under noise conditions.
@@ 2.2 Speech Separation
The proposed SeamlessEdit framework leverages StoRM [18] as a speech separation module to efficiently separate human voice and background noise in a few diffusion steps. This separation module integrates predictive losses with diffusion processes, allowing the reduction of complexity and suppression of artifacts. In contrast to the costly iterative refinement process in other speech separation methods [19, 20], the proposed model is efficient and robust to distinct acoustic conditions. The speech separation operation can be expressed as $X_{s},X_{n}=\text{StoRM}(X)$ where $X_{s}$ is the separated speech signal and $X_{n}$ is the background noise. To achieve a good result for speech separation, $X_{s}$ should preserve the low- and high-frequency speech harmonics even if $X$ interferes with audible noise patterns.
@@ 2.3 Frequency-Band Aware Noise Suppression
Speech signal separation is challenging if its support in the frequency domain is intertwined with that of the noise, which can bias the predictive posterior estimation in StoRM for denoising scoring. Therefore, we propose a noise suppression module to strengthen speech intelligibility while attenuating residual noise for unbiased voice representation. We employ sparse Bayesian learning (SBL) [21, 22] to estimate the sparse frequency bands of the human voice and thus eliminate the residual noise whose spectrum is usually uniformly distributed. SBL assumes that signals exhibit sparsity in an appropriate transform domain.
Specifically, $X_{s}$ is decomposed into the human voice $X_{p}$ and the residual noise $N$ as follows:
@@
where $\gamma_{i}$ controls sparsity, $\lambda$ is a regularization parameter balancing the sparsity and the reconstruction accuracy, and $\epsilon$ prevents division by zero. The iteration terminates when $\gamma$ converges beyond a small threshold $\delta$: $|\gamma^{(t+1)}-\gamma^{(t)}|<\delta$.
Once the sparse coefficients $X_{p}$ are estimated, the speech signal undergoes a Butterworth filtering to enhance the desired frequency components while suppressing residual noise. The transfer function of the Butterworth filter is defined as:
@@ 2.4 Neural Codec Editing Model for TTS
We adopt VoiceCraft [14] as the neural codec editing model for TTS to perform insertion, deletion, and substitution flexibly and efficiently. Its causal masking and delayed stacking mechanisms enhance the Encodec backbone for discrete tokenization while simplifying diffusion steps.
For noisy speech editing, we use a separated speech input $X_{s}=\{X_{s,1},…,X_{s,T}\}$ and the target region mask $M=\{m,…,m+k\}$ within the neural codec editing model $f$ to generate an edited speech $X_{e}=\{X_{e,1},…,X_{e,T}\}$. The edited region $\{X_{e,m},…,X_{e,m+k}\}$ maintains quality consistent with non-edited regions, ensuring seamless integration. VoiceCraft is adopted to perform autoregressive sequence prediction conditioned on the speech transcript.
To further enhance speech quality, we apply SBL filtering to suppress residual noise and refine the spectral characteristics of the separated speech. The filtered output $X_{l}=\{X_{l,1},…,X_{l,T}\}$ is served as the input of the TTS editing model $g$ to generate the edited speech $X_{le}=\{X_{le,1},…,X_{le,T}\}$. Given a masked target region $M_{l}$, the TTS editing model modifies this segment while preserving the surrounding regions. The autoregressive generation process, conditioned on linguistic and acoustic representations, ensures natural and coherent synthesis.
@@ 2.5 In-Context Refinement
The auditory quality of the generated speech from the neural codec editing model $f$ described in §$\S$2.4 with input $X_{s}$ is naturally limited by its intricate conjunction between the human voice and background noise. Inspired by the success of prompt design [23, 24] and in-context learning [25] in speech recognition, we introduce the in-context learning strategy to refine the editing process by integrating the low-frequency speech $X_{le}$ described in §$\S$2.4 to emphasize frequency components of human voices. It should be noted that previous in-context learning studies for speech generation usually enhance descriptive attributes such as timbre or spoken content [26, 27] without incorporating additional acoustic information to refine the speech quality.
The proposed in-context refinement strategy is implemented by a MultiHead attention module before the TTS model $f$ that applies low-frequency embedding $X_{le}$ to enhance $X_{s}$ in §$\S$2.2 with intrinsic human voice characteristics. In Figure 1, the input of $Q$ is embedded with $X_{s}$ while the inputs of $K$ and $V$ are embedded with low-frequency speech $X_{l}$. $Q$, $K$, and $V$ embeddings are derived by projecting with learnable projection matrices $W_{i}^{Q}X_{s}$, $W_{i}^{K}X_{l}$, and $W_{i}^{V}X_{l}$. Then, the attention of the $i^{th}$ head is determined as follows:
@@ 3 Experiments
Figure 2: Mel-spectrogram case study of (a) clean and (e) noisy ground truth. Different editing stages include (b) separated speech $X_{s}$, (c) noise-suppressed speech $X_{l}$, (d) SeamlessEdit processed clean speech $X_{e}$, and (f) the final noisy editing result $Y$ of the proposed SeamlessEdit model. Red texts and boxes denote the edited regions.
@@ 3.1 Evaluation Setup
Our experiment is conducted upon the noisy test set of the EARS-WHAM dataset for three editing tasks, including insertion, short replacement, and long replacement. Insertion and short replacement are performed on 1-6 words editing and long replacement is conducted on 7-12 words. The compared baselines include FluentSpeech and Voicecraft. Although FluentSpeech cannot perform insertion, we regard it for the short and long replacement tasks. We also apply the ground truth and Voicecraft’s clean condition results as the upper-bound reference.
We adopt the evaluation metrics of the word error rate (WER), the perceptual edited score (PES), the naturalness mean opinion score (NMOS), and the subjective mean opinion score (SMOS). The WER measures transcription intelligibility by the Whisper medium [28]. The PES indicates the level of perceived edited cues and its value is within the range of [0, 2]. The NMOS assesses relative speech naturalness through pairwise comparisons of fluency and coherence. The SMOS provides an absolute quality rating based on human perception. We recruited 30 subjects to perform the subjective evaluation for the NMOS and the SMOS whose values were reported within the range of [1, 5]. A lower WER, a lower PES, a higher NMOS, and a higher SMOS mean better performance.
Figure 3: Spectral centroid and bandwidth statistics for short replacement results. We present the result of each editing step in Figure 2 for clean and noisy conditions.
Implementation Details: A single NVIDIA A16 GPU is applied for implementation. The TTS model described in §$\S$2.4 is implemented following the setting in Voicecraft using a tokenizer with the frame rate of 50kHz and the vocabulary size of 2048. The main architecture comprises 16 transformer layers each with 2048 hidden nodes and 8192 output nodes, followed by 4 dense layers for logit prediction. The hyperparameters for noise suppression described in §$\S$2.3 are $a=[1,0,0.486029,0,0.017665]$ and $b=[0.093981,-0.375923,0.563885,-0.375923,0.093981]$.
@@ 3.2 Main Results
As shown in Table 1, the proposed SeamlessEdit model achieves 3.78, 3.56, and 3.65 NMOS in insertion, short, and long replacement tasks, respectively. The SMOS results of SeamlessEdit also outperform those of other methods. The performance of the existing method is significantly degraded in noisy scenarios. The state-of-the-art model, VoiceCraft, shows -1.30, -1.35, and -1.39 NMOS differences in insertion, short, and long replacement. Similarly, the SMOS dropped by an average of 1.39 for the three tasks. This is because the noise hinders the neural codec model from capturing real human voice characteristics. FluentSpeech exhibits a comparable WER to VoiceCraft but a lower MOS in replacement tasks due to its direct segment replacement, which disrupts waveform continuity.
Moreover, the proposed SeamlessEdit model attains 0.72 PES, demonstrating that the proposed model has smoother and more natural editing outcomes. We examine the results of SeamlessEdit w/o in-context learning (ICL) to illuminate the effects of the in-context refinement. As a result, SeamlessEdit w/o ICL obtains a comparable SMOS with a lower NMOS and a higher PES compared to SeamlessEdit. The results verify that the refinement mechanism can suppress the noise effect and improve the speech edit performance.
@@ 3.3 Spectrogram Analysis
Figure 2 and Figure 3 show the visual comparison of Mel spectrograms and quantitative spectral characteristics, highlighting the effects of different processing steps. In Figure 2, we compare the edited speech in each stage to the clean ground truth in Figure 2 (a). The separated speech $X_{s}$ in Figure 2 (b) shows indistinct and blurry boundaries around the voiced segments owing to the residual noises. The SBL-filtered spectrogram in Figure 2 (c) has suppressed more noise but excessive filtering can over-attenuate high-frequency components, leading to a muffled sound. SeamlessEdit in Figure 2 (d) balances the goals of noise removal and the preservation of high-frequency voice component via ICL refinement. Therefore, the final noisy editing results in Figure 2 (f) are indistinguishable from the noisy ground truth in Figure 2 (e). There is a challenge of distinctly isolating female voices from background noise, as Sample 2 shows. The residual frequency components still appear in the spectrogram, despite applying separation and noise suppression techniques.
We leverage the spectral centroid and bandwidth [29, 30] to measure the intensity-weighted average frequency and the average frequency band disparity in Figure 3. These two metrics have higher values for the separated speech $X_{s}$ and lower values for the noise-suppressed speech $X_{l}$. The spectral bandwidth of $X_{l}$ stays close to that of clean speech while the centroid values are lower. The results are consistent with Figure 2 that $X_{l}$ invades parts of high-frequency voices and $X_{s}$ remains intertwined with noise. The difficulty of resolving the overlapped high-frequency bands manifests in female speakers. In Figure 3 (c), the female voices are more susceptible to noise interference. Similar evidence is shown in the spectral bandwidth of $X_{e}$ that the average value of the females is higher than that of the males. Intriguingly, increased timbre preservation can be achieved by the proposed ICL refinement mechanism that guides the synthesis with the noise-suppressed signal as a reference example.

### v2

@@ Abstract
With the fast development of zero-shot text-to-speech technologies, it is possible to generate high-quality speech signals that are indistinguishable from the real ones. Speech editing, including speech insertion and replacement, appeals to researchers due to its potential applications. However, existing studies only considered clean speech scenarios. In real-world applications, the existence of environmental noise could significantly degrade the quality of generation. In this study, we propose a noise-resilient speech editing framework, SeamlessEdit, for noisy speech editing. SeamlessEdit adopts a frequency-band-aware noise suppression module and an in-content refinement strategy. It can well address the scenario where the frequency bands of voice and background noise are not separated. The proposed SeamlessEdit framework outperforms state-of-the-art approaches in multiple quantitative and qualitative evaluations.
Index Terms—  speech editing, noisy speech, in-context learning, speech synthesis
@@ 1 Introduction
Several techniques can effectively improve the quality of audio generation and eliminate the acoustic discrepancy in the segment boundaries. Typically, fade-in and fade-out techniques are used to smooth the transitions between cut-and-paste audio segments [4]. However, intricate editing procedures may generate artifacts and degrade audio quality. Text prompts are viable for generating the speech signal following a specific style  [5]. P-Flow in  [6] uses a prompt-based speaker-conditioned text representation to improve quality. In addition, context-aware methods have been proposed to eliminate unnatural edited boundaries  [7]. FluentSpeech [8] alleviates the over-smoothing problem using iterative refinement for stutter removal. Mapache [9] addresses the problem of left-to-right generation in autoregressive (AR) models by a non-AR parallel transformer. FluentEdit [10] applies the acoustic and prosody consistency loss when training the diffusion model for text-based editing. InstructSpeech [11] is a multitask model and applies a large language model for training.
@@
This study presents SeamlessEdit, a noise-robust speech editing framework that introduces in-context refinement for natural integration of speech and background noise. Unlike prior methods that assume clean inputs, SeamlessEdit is designed for real-world acoustic conditions. The framework first separates speech and ambient noise with a score-based diffusion model, then applies sparse Bayesian learning (SBL) filtering to preserve key speech structures while reducing residual noise. The refined signal is incorporated into the decoder for in-context refinement, enabling seamless boundary transitions and enhanced perceptual quality.
Experiments on the EARS-WHAM dataset [16] demonstrate significant improvements, achieving 29.01% NMOS and 24.28% SMOS gains over state-of-the-art systems. Human evaluations further confirm that SeamlessEdit produces audio nearly indistinguishable from natural recordings, effectively blending speech, edits, and ambient noise. Case studies also highlight robustness in challenging scenarios such as overlapping talkers, spontaneous pauses, and dynamic noise. Together, these results establish SeamlessEdit as a practical solution for applications including podcast editing, interview restoration, and archival enhancement.
Fig. 1: The proposed SeamlessEdit framework separates human voice $X_{s}$ and suppresses residual noise to derive an edited speech $X_{le}$. An in-context refinement enhances the editing of $X_{s}$ using $X_{le}$ for indistinguishable noisy editing results. The SBL filter is also adopted to improve the robustness to noise.
@@ 2 Methods
For our experiments, we used the EARS-WHAM dataset, which combines high-quality 16kHz speech from the EARS corpus [16] with diverse background noises from WHAM! [17]. The dataset provides expressive and conversational speech mixed with real-world noise at 198 SNR levels. We performed evaluation on both clean and noisy subsets, using 886 speech samples paired with 594 noise recordings, with an average speech duration of 15.6 seconds.
@@ 2.1 SeamlessEdit: Noisy Speech Editing Framework
Given an input speech sequence $X=\{X_{1},\ldots,X_{T}\}$ of length $T$, the $m^{\text{th}}$ to $(m+k)^{\text{th}}$ samples specify the target editing region. The edited speech is denoted as
$X_{e}=\{X_{1},\ldots,X_{m},\ldots,X_{m+k},\ldots,X_{T}\}.$
The proposed SeamlessEdit framework consists of a speech separation module (Section 2.2) followed by frequency-aware noise suppression (Section 2.3), which extracts the separated signal $X_{s}$, the noise-suppressed signal $X_{l}$, and the background signal $X_{n}$. An in-context refinement strategy further enhances the pre-trained editing model (Section 2.4) to generate edited voices under noisy conditions.
@@ 2.2 Speech Separation
The proposed SeamlessEdit framework leverages StoRM [18] as a speech separation module to efficiently separate human voice and background noise in a few diffusion steps. This separation module integrates predictive losses with diffusion processes, allowing the reduction of complexity and suppression of artifacts. In contrast to the costly iterative refinement process in other methods, the proposed speech separation model is efficient and robust to distinct acoustic conditions. The speech separation operation can be expressed as $X_{s},X_{n}=\text{StoRM}(X)$ where $X_{s}$ is the separated speech signal and $X_{n}$ is the background noise. To achieve a good result for speech separation, $X_{s}$ should preserve low- and high-frequency speech harmonics even if $X$ is interfered with audible noise patterns.
@@ 2.3 Frequency-Band Aware Noise Suppression
Speech signal separation is challenging if its support in the frequency domain is intertwined with that of the noise, which can bias the posterior estimation result. Therefore, we propose a noise suppression module to strengthen speech intelligibility while attenuating residual noise for unbiased voice representation. We employ sparse Bayesian learning (SBL) [19, 20] to estimate the sparse frequency bands of the human voice and thus eliminate residual noise whose spectrum is usually uniformly distributed. SBL assumes that signals exhibit sparsity in an appropriate transform domain.
Specifically, $X_{s}$ is decomposed into the human voice $X_{p}$ and residual noise $N$ as follows:
@@
where $\gamma_{i}$ controls sparsity, $\lambda$ is a regularization parameter to balance the sparsity and the reconstruction accuracy, and $\epsilon$ prevents the division from being zero. The iteration ends when $\gamma$ converges beyond a small threshold $\delta$: $|\gamma^{(t+1)}-\gamma^{(t)}|<\delta$.
Once the sparse coefficients $X_{p}$ are estimated, the speech signal undergoes a Butterworth filter to enhance the desired frequency components while suppressing residual noise. The transfer function of the Butterworth filter is defined as follows:
@@ 2.4 Neural Codec Editing Model for TTS
We adopt VoiceCraft [14] as the neural codec editing model for TTS to perform insertion, deletion, and substitution flexibly and efficiently. Its causal masking and delayed stacking mechanisms enhance the codec backbone for discrete tokenization while simplifying diffusion steps.
For noisy speech editing, we use a separated speech input $X_{s}=\{X_{s,1},…,X_{s,T}\}$ and the target region mask $M=\{m,…,m+k\}$ within the neural codec editing model $f$ to generate edited speech signals $X_{e}=\{X_{e,1},…,X_{e,T}\}$. The edited region $\{X_{e,m},…,X_{e,m+k}\}$ maintains quality consistent with non-edited regions, ensuring seamless integration. VoiceCraft is adopted to perform autoregressive sequence prediction conditioned on the speech transcript.
To further enhance speech quality, we apply SBL filters to suppress residual noise and refine the spectral characteristics of separated speech signals. The filtered outputs $X_{l}=\{X_{l,1},…,X_{l,T}\}$ are served as the inputs of the TTS editing model $g$ to generate edited speech signals $X_{le}=\{X_{le,1},…,X_{le,T}\}$. Given a masked target region $M_{l}$, the TTS editing model modifies this segment while preserving the surrounding regions. The autoregressive generation process, conditioned on linguistic and acoustic representations, ensures natural and coherent synthesis.
@@ 2.5 In-Context Refinement
In $\S$2.4, the auditory quality of the generated speech signal from the neural codec editing model $f$ is naturally limited by its intricate conjunction between the human voice and background noise. Inspired by the success of prompt design [21, 22] and in-context learning [23] in speech recognition, we introduce an in-context learning strategy to refine the editing process by integrating the low-frequency speech $X_{le}$ described in $\S$2.4 to emphasize the frequency components of human voices. Note that previous in-context learning studies for speech generation usually enhance descriptive attributes such as timbre or spoken content [24, 25] without incorporating additional acoustic information to refine the speech quality.
The proposed in-context refinement strategy is implemented by a MultiHead attention module before the TTS model $f$ that applies low-frequency embedding $X_{le}$ to enhance $X_{s}$ in $\S$2.2 with intrinsic human voice characteristics. In Figure 1, the input of $Q$ is embedded with $X_{s}$ while the inputs of $K$ and $V$ are embedded with the low-frequency speech signals $X_{l}$. $Q$, $K$, and $V$ embeddings are derived by projecting with learnable projection matrices $W_{i}^{Q}X_{s}$, $W_{i}^{K}X_{l}$, and $W_{i}^{V}X_{l}$, respectively. Then, the attention of the $i^{th}$ head is determined as follows:
@@ 3 Experiments
Fig. 2: Mel-spectrograms of each stage of the proposed model; (a) the clean signal; (e) the noisy ground truth. Different editing stages include (b) separated speech $X_{s}$, (c) noise-suppressed speech $X_{l}$, (d) SeamlessEdit processed clean speech $X_{e}$, and (f) the final noisy editing result $Y$ of the proposed SeamlessEdit model. Red texts and boxes denote the edited regions.
@@ 3.1 Evaluation Setup
Experiments are conducted on the noisy test set of the EARS-WHAM dataset for three editing tasks, including insertion, short replacement, and long replacement. Insertion and short replacement are performed on 1 6-word editing, and long replacement is conducted on 7 12-word editing. The compared baselines include FluentSpeech and Voicecraft. Although FluentSpeech cannot perform insertion, we regard it for short and long replacement tasks. We also apply the ground truth and Voicecraft’s clean condition results as the upper-bound reference.
We adopt the evaluation metrics of the word error rate (WER), the perceptual edited score (PES), the naturalness mean opinion score (NMOS), and the subjective mean opinion score (SMOS). The WER measures the transcription intelligibility by the Whisper medium [26]. The PES indicates the level of perceived edited cues and its value is within [0, 2]. The NMOS assesses relative speech naturalness through pairwise comparisons of fluency and coherence. The SMOS evaluates absolute quality based on human perception. We recruited 30 subjects to perform subjective evaluation for the NMOS and the SMOS whose values were reported within the range of [1, 5]. A lower WER, a lower PES, a higher NMOS, and a higher SMOS mean better performance.
Fig. 3: Spectral centroid and bandwidth statistics for short replacement results. We present the result of each editing step in Figure 2 for clean and noisy conditions. An ideal editing result should be indistinguishable from the noisy ground.
Implementation Details: A single NVIDIA A16 GPU is applied for implementation. The TTS model described in $\S$2.4 is implemented following the setting in Voicecraft using a tokenizer with the frame rate of 50kHz and the vocabulary size of 2048. The main architecture comprises 16 transformer layers (each has 2048 hidden nodes and 8192 output nodes) followed by 4 dense layers for logit prediction. The hyperparameters for noise suppression described in $\S$2.3 are $a=[1,0,0.486029,0,0.017665]$ and b=[0.093981,0.375923,0.563885,0.375923,0.093981]b=[0.093981,-0.375923,0.563885,-0.375923,0.093981].
@@ 3.2 Main Results
As shown in Table 1, the proposed SeamlessEdit model achieves 3.78, 3.56, and 3.65 NMOS in insertion, short, and long replacement tasks, respectively. The SMOS results of SeamlessEdit also outperform those of other methods. The performance of the existing method is significantly degraded in noisy scenarios. The state-of-the-art model, VoiceCraft, shows -1.30, -1.35, and -1.39 NMOS differences in insertion, short, and long replacement. Similarly, the SMOS dropped by an average of 1.39 for the three tasks. This is because noise hinders the neural codec model from capturing real human voice characteristics. FluentSpeech exhibits a comparable WER to VoiceCraft but a lower MOS in replacement tasks due to its direct segment replacement, which disrupts waveform continuity.
Moreover, the proposed SeamlessEdit model attains 0.72 PES, demonstrating its smoother and more natural editing results. We examine the results of SeamlessEdit w/o in-context learning (ICL) to illuminate the effects of in-context refinement. As a result, SeamlessEdit w/o ICL obtains a comparable SMOS, a lower NMOS, and a higher PES compared to SeamlessEdit. The results verify that the refinement mechanism can suppress the noise effect and improve the speech editing performance.
@@ 3.3 Spectrogram Analysis
Fig. 2 and Fig. 3 show the visual comparison of Mel spectrograms and quantitative spectral characteristics, highlighting the effects of different processing steps. In Fig. 2, we compare the edited speech in each stage with the clean ground truth in Fig. 2 (a). The separated speech $X_{s}$ in Fig. 2 (b) shows indistinct and blurry boundaries around the voiced segments owing to residual noise. The SBL-filtered spectrogram in Fig. 2 (c) has suppressed more noise, but excessive filtering can over-attenuate high-frequency components, leading to a muffled sound. SeamlessEdit in Fig. 2 (d) balances the goals of noise removal and the preservation of high-frequency voice component via ICL refinement. Therefore, the final noisy editing results in Fig. 2 (f) are indistinguishable from the noisy ground truth in Fig. 2 (e). There is a challenge of distinctly isolating female voices from background noise, as Sample 2 shows. The residual frequency components still appear in the spectrogram despite applying separation and noise suppression techniques.
We used the spectral centroid and the bandwidth [27, 28] to measure the intensity-weighted average frequency and the average frequency band disparity in Fig. 3. These two metrics have higher values for the separated speech signal $X_{s}$ and lower values for the noise-suppressed signal $X_{l}$. The spectral bandwidth of $X_{l}$ stays close to that of the clean speech signal while the centroid values are lower. The results are consistent with Fig. 2 that $X_{l}$ invades parts of high-frequency voices and $X_{s}$ remains intertwined with noise. The difficulty of resolving the overlapped high-frequency bands manifests in female speakers. In Fig. 3, the female voices are more susceptible to noise interference. Similar evidence is shown in the spectral bandwidth of $X_{e}$ that the average value of females is higher than that of males. Intriguingly, increased timbre preservation can be achieved by the proposed ICL refinement mechanism that leverages the synthesis of the noise-suppressed signal as a reference example.